# MalMixer: Few-Shot Malware Classification with Retrieval-Augmented Semi-Supervised Learning

### v2

@@ III-B Retrieval-Based Augmentation Pipeline
We first divide static features into interpolatable (denoted $\mathcal{I}$) and non-interpolatable (denoted $\mathcal{N}$). The designation of interpolatable vs. non-interpolatable follows an intuitive rule: a feature is interpolatable if and only if that feature is numerical in nature and perturbing that feature still results in sensible values. For example, the byte entropy of a file is interpolatable since any interpolated value remains a plausible entropy for some executable. Conversely, a hash of an imported filename is non-interpolatable, as not all hashes correspond to valid filenames. As such, typical interpolatable features include byte entropy histograms, Shannon’s entropy of printable characters, and the occurrence count of string literals. Typical non-interpolatable features encompass major OS versions and hash values of libraries / section names. The rule for splitting features into interpolatale vs non-interpolatable does not change based on features: for example, if using control-flow-graph features, those features non-numerical in nature (e.g. control dependencies) will be non-interpolatable.
With domain knowledge and research effort, the designation of interpolatable vs non-interpolatable features can be performed quickly on any arbitrary malware feature representations: for instance, in our studies, we processed a 2381-dimensional static feature representation of malware within five hours by an author with a security background. The set of features used in this study is the EMBER raw features [42], which are used in the BODMAS and MOTIF datasets introduced in Section V-A. The particular interpolatable vs. non-interpolatable split that we designated for our experiments can be found in Section F in the Appendix.
Regarding the impact of the interpolatable vs. non-interpolatable split, a way of not following such split would be interpolating all features in the same way, which demonstrated degraded performance (Section VII-E).
@@
Next, we project malware features into two embeddings: (1) an invariant embedding, capturing the commonalities between interpolatable and non-interpolatable features from the same malware sample; and (2) a dissimilar embedding, emphasizing the unique aspects that differentiate interpolatable from non-interpolatable features. We use these embeddings to capture interdependent features and family-common features that are used to guide the augmentation in steps (iii) and (iv).
@@
To achieve this projection into two embedding spaces, we design a separate encoder-decoder structure for each of $s_{i}$ and $s_{n}$. Then, we perform feature reconstruction using ${s_{i}}^{\prime}=\psi_{i}(\phi_{i}(s_{i})),~{}{s_{n}}^{\prime}=\psi_{n}(\phi_{n}(s_{n})),$ where $\phi_{i}$ (and $\phi_{n}$) and $\psi_{i}$ (and $\psi_{n}$) represent the encoders and decoders for $s_{i}$ (and $s_{n}$). The reconstructed variables are denoted as ${s_{i}}^{\prime}$ and ${s_{n}}^{\prime}$. The corresponding loss for malware feature reconstruction is $\mathcal{L_{R}}=\frac{1}{N}\sum_{k=1}^{N}\left(||s_{i}-{s_{i}}^{\prime}||^{2}_{2}+||s_{n}-{s_{n}}^{\prime}||^{2}_{2}\right)_{k},$ where $\mathcal{L}_{r}$ represents the reconstruction loss, and $N$ denotes the total number of data points in a batch. By leveraging the encoder-decoder architecture, developers can transform static features of malware into a more compact representation that captures the underlying structure and semantics of the executable.
@@
To separate domain-specific features from domain-invariant features during the reconstruction process, we introduce two auxiliary tasks for the encoder-decoder. As shown in Figure 3, we divide the hidden features $h_{i}\coloneqq\phi_{i}(x_{i})$ and $h_{n}\coloneqq\phi_{n}(x_{n})$ evenly, splitting them into two parts each. For instance, if the original dimension of $h_{i}$ is 512, each resulting vector will have a dimension of 256. We then perform similarity and dissimilarity learning on these vectors. Specifically, we define the two parts of the split vector as the similar vectors, denoted as $h_{i}^{sim}$ and $h_{n}^{sim}$, and the dissimilar vectors, denoted as $h_{i}^{dis}$ and $h_{n}^{dis}$, obtaining $h_{i}=h_{i}^{sim}\oplus h_{i}^{dis}$ and $h_{n}=h_{n}^{sim}\oplus h_{n}^{dis}$. We enforce the dissimilar vectors to be distant from one another, while ensuring the similar vectors remain close together. This is expressed as $\mathcal{L_{S}}=\frac{1}{N}\sum_{k=1}^{N}||h_{i}^{sim}-h_{n}^{sim}||_{2}^{2},~{}\mathcal{L_{D}}=\frac{1}{N}\sum_{k=1}^{N}\left(\max\left(0,5-||h_{i}^{dis}-h_{n}^{dis}||_{2}^{2}\right)\right)_{k},$ where $\mathcal{L_{S}}$ and $\mathcal{L_{D}}$ denote similarity loss and dissimilarity loss, respectively. The total loss is $\mathcal{L_{R}}+\mathcal{L_{D}}+\mathcal{L_{S}}$.
@@
After training the encoder-decoder model, we lock the parameters of both encoders $\phi_{i}$ and $\phi_{n}$. Then, we apply $\phi_{i}$ to $\mathcal{S}_{i}$, $\phi_{n}$ to $\mathcal{S}_{n}$, and obtain an embedding of interpolatable features $\mathcal{H}_{i}:=\phi_{i}(\mathcal{S}_{i})$ and an embedding of non-interpolatable features $\mathcal{H}_{n}:=\phi_{n}(\mathcal{S}_{n})$. Note that $\mathcal{H}_{i}=concat(\mathcal{H}_{i}^{sim},\mathcal{H}_{i}^{dis})$, where $\mathcal{H}_{i}^{sim}$ is the invariant embedding of interpolatable features containing all $h_{i}^{sim}$, and $\mathcal{H}_{i}^{dis}$ is dissimilar embedding of interpolatable features containing all $h_{i}^{dis}$. $concat$ is the pairwise feature concatenation of $(h_{i}^{sim},h_{i}^{dis})$ such that $h_{i}^{sim},h_{i}^{dis}$ belong to the same $h_{i}$. Similarly, $\mathcal{H}_{n}=concat(\mathcal{H}_{n}^{sim},\mathcal{H}_{n}^{dis})$.
@@
As shown in Figure 4, we refer to $\mathcal{H}_{n}$ as the N-retrieval embedding, and $\mathcal{H}_{i}^{sim}\cup\mathcal{U}_{n}^{sim}$, which is the invariant embedding of both interpolatable and non-interpolatable features, as I-alignment embedding. Note that a one-to-one correspondence exists between $\mathcal{S}$ (the original malware feature representation set), the N-retrieval embedding (representing non-interpolatable features), and the I-alignment embedding (representing common aspects between interpolatable and non-interpolatable features). For every original malware sample in $\mathcal{S}$, there is a projection of its non-interpolatable features in the N-retrieval embedding. Similarly, the shared aspects between each sample’s interpolatable and non-interpolatable features are mapped onto the I-alignment embedding.
@@
Once the N-retrieval and I-alignment databases are obtained, we proceed to steps (iii) and (iv), which mutate a malware instance by mixing it with a similar instance, as motivated by our insights outlined in Section I. Steps (iii) and (iv) are illustrated by Figure 5.
@@
Figure 6: Illustration of similarities shared among families in BODMAS. The left subfigure displays such similarities in the original static feature space $\mathcal{S}$. The middle subfigure depicts shared similarities in the N-Retreival embedding $\mathcal{H}_{n}$. The right subfigure represents similarities between interpolatable and non-interpolatable features of similar malware when projected to the I-Alignment invariance embedding $\mathcal{H}_{i}^{sim}\cup\mathcal{H}_{n}^{sim}$.
@@
For task (a), we identify the top-k similar malware samples in $\mathcal{S}$ compared to $s$ through constructing a k-nearest-neighbor graph. We select L2 distance as the similarity metric for the $k$-NN graph. From the top-k similar malware samples, we randomly select one neighbor, denoted $s^{\prime}$.
@@
where $\alpha$ is the same $\alpha$ used above. We consequently select the top-k nearest neighbors of $\tilde{h}_{n}$ in $H_{n}$ using L2 distance as similarity metric. The selected neighbors, denoted $\{\hat{h}_{n_{1}},\cdots,\hat{h}_{n_{k}}\}$, are regarded as candidates in the N-Retrieval embedding space representing a mixture of non-interpolatable features of $s$ and $s^{\prime}$. A more detailed motivation for performing task (c) and the following step (iv) in the embedding spaces, rather than of the original space, will be provided in Section III-C.
@@
In doing so, we choose one set of non-interporlatable features that aligns most closely with the mixed interpolatable features in the I-alignment embedding. Recall that $\hat{h}_{n_{j}}$ has a one-on-one correspondence with some $\tilde{s}_{n}\in\mathcal{S}_{n}$ in the original feature space. As a result, we use $\tilde{s}_{n}$ as the mixture of $s$ and $s^{\prime}$ by their non-interpolatable features. We define our augmentation method $\mathcal{A}$ on $s\in\mathcal{S}$ such that $\mathcal{A}(s)=\tilde{s}_{i}\oplus\tilde{s}_{n}$.
@@
In essence, $\mathcal{A}(s)$ is a concatenation of a direct mutation on the interpolatable features of $s$ and a real set of non-interpolatable features retrieved from another training sample. Steps (iii) and (iv) address the problem that arises when mixing non-interpolatable features from two samples, which typically results in a semantically nonsensical mixture. This issue is mitigated by selecting non-interpolatable features from real samples that are akin to the mixed instance (step (iii)) and compatible with the mutated interpolatale features (step (iv)).
@@
During the augmentation process, establishing a one-on-one between space $\mathcal{S}$ and the hidden spaces $\mathcal{H}_{i},\mathcal{H}_{n}$ allows the model to optimize across both the original interpolatable feature space and the non-interpolatable hidden space. This results in a more realistic synthetic sample. We provide a detailed justification for our design choices in the next subection and a intuitive walkthrough of our proposed augmentation pipeline in Section IV. The augmentation method is formulated in Algorithm 1, provided in the Appendix.
III-C Assumptions and Insights
We make three key assumptions that underpin our augmentation approach based the insights described in Section I. In this subsection, we describe and validate these assumptions.
(i) We assume that same-family malware are similar in the original static feature space ($\mathcal{S}$). Consequently, we mutate each malware sample $s$ by mixing it with one of its top-k neighbors, presuming they are instances of the same family and thus semantically similar. We hypothesize that mixing two such samples likely yields new sample with more realistic features than existing augmentation techniques.
(ii) We assume that non-interpolatable features from similar malware are projected closer together within an embedding space. Thus, we expect that mixing these features’ projections in N-Retrieval embedding $\mathcal{H}_{n}$ via linear combination will neighbor other non-interpolatable features corresponding to similar functionality.
(iii) We assume that using alignment techniques, interpolatable and non-interpolatable features from the same malware are projected onto similar subspaces within an embedding space. Thus, similarity within the I-Alignment invariance embedding $\mathcal{H}_{i}^{sim}\cup\mathcal{U}_{n}^{sim}$ indicates compatibility between a pair of interpolatable and non-interpolatable features (i.e., some pairs of features may be interdependent). Accordingly, we use similarity search between interpolatable and non-interpolatable features in this embedding space to align them and maximize feature consistency within samples we synthesize.
Similarity analysis Figure 6 visualizes these insights using T-SNE [45] over samples in the BODMAS dataset. We measure similarity using Euclidean distance [45]. Assumption (i) is supported by the organized clusters of malware families in Figure 6-left. Numerically, we also verified that, within BODMAS-20, each malware sample shares the same family as $95.86\%$ of its top-5 nearest neighbors. For assumption (ii), Figure 6-middle shows that non-interpolatable features form family-based clusters in the N-Retrieval embedding, $\mathcal{H}_{n}$. In contrast, most non-interpolatable features do not form clusters in the original feature space $\mathcal{S}_{n}$. Thus, we mix non-interpolatable features in the embedding space $\mathcal{H}_{n}$ (Equation (4)). Finally, assumption (iii) is supported by the overlapping subspaces of same-family features in the invariance embedding $\mathcal{H}_{i}^{sim}\cup\mathcal{H}_{n}^{sim}$, as illustrated in Figure 6-right. Consequently, similarity between projected interpolatable and non-interpolatable features indicates higher compatibility.
@@ III-C Assumptions and Insights
Figure 7: Case study workflow. In the left subfigure, we identify several non-interpolatable features that represent a blend of the two chosen samples in the embedding space. In the middle subfigure, we select the non-interpolatable features that best aligns with the combined interpolatable features in the embedding space. In the right subfigure, we map the synthetic malware, which consists of the chosen mixed features, back to the original feature space.
IV Motivating Example

### v4

@@ 3.2 Retrieval-Based Data Augmentation Pipeline
We first divide static features into interpolatable (denoted $\mathcal{I}$) and non-interpolatable (denoted $\mathcal{N}$). The designation of interpolatable vs. non-interpolatable follows an intuitive rule: a feature is interpolatable if and only if that feature is numerical in nature and perturbing that feature still results in sensible values. For example, the byte entropy of a file is interpolatable since any interpolated value remains a plausible entropy for some executable. Conversely, a hash of an imported filename is non-interpolatable, as not all hashes correspond to valid filenames. As such, typical interpolatable features include byte entropy histograms, Shannon’s entropy of printable characters, and the occurrence count of string literals. Typical non-interpolatable features encompass major OS versions and hash values of libraries / section names. The rule for splitting features into interpolatable vs non-interpolatable does not change based on features: for example, if using control-flow-graph features, those features non-numerical in nature (e.g., control dependencies) will be non-interpolatable.
With domain knowledge and research effort, the designation of interpolatable vs non-interpolatable features can be performed quickly on any arbitrary malware feature representations: for instance, in our studies, we processed a 2381-dimensional static feature representation of malware within five hours by an author with a security background. The set of features used in this study is the EMBER raw features [42], which are used in the BODMAS and MOTIF datasets introduced in Section 5.1. The particular interpolatable vs. non-interpolatable split that we used for our experiments is described in Appendix G.
We note that, when interpolating all features in the same way (i.e., without splitting them), classification performance degraded (Section 7.5).
@@
Parts (ii)–(iv) below present the mathematical formulations of the next steps in our proposed pipeline. However, readers may find it helpful to first refer to motivations behind our design in Section 3.3 and/or a case study in Section 4 that intuitively walks through steps (i)–(iv).
@@
Next, we project malware features into two embeddings: (1) an invariant embedding, capturing the commonalities between interpolatable and non-interpolatable features from the same malware sample; and (2) a dissimilar embedding, emphasizing the unique aspects that differentiate interpolatable from non-interpolatable features. We use these embeddings to capture interdependent features and family-common features that are used to guide the data augmentation in steps (iii) and (iv).
@@
To achieve this projection into two embedding spaces, we design a separate encoder-decoder structure for each of $s_{i}$ and $s_{n}$. Then, we perform feature reconstruction using ${s_{i}}^{\prime}=\psi_{i}(\phi_{i}(s_{i})),~{}{s_{n}}^{\prime}=\psi_{n}(\phi_{n}(s_{n})),$ where $\phi_{i}$ (and $\phi_{n}$) and $\psi_{i}$ (and $\psi_{n}$) represent the encoders and decoders for $s_{i}$ (and $s_{n}$). The reconstructed variables are denoted as ${s_{i}}^{\prime}$ and ${s_{n}}^{\prime}$. The corresponding loss for malware feature reconstruction is $\mathcal{L_{R}}=\frac{1}{N}\sum_{k=1}^{N}\left(||s_{i}-{s_{i}}^{\prime}||^{2}_{2}+||s_{n}-{s_{n}}^{\prime}||^{2}_{2}\right)_{k},$ where $\mathcal{L}_{\mathcal{R}}$ represents the reconstruction loss, and $N$ denotes the total number of data points in a batch. The encoder-decoder architecture enables transforming static features of malware into a more compact representation capturing the underlying structure and semantics of the executable.
@@
To separate domain-specific features from domain-invariant features during the reconstruction process, we introduce two auxiliary tasks for the encoder-decoder. As shown in Figure 3, we divide the encoded hidden features $h_{i}\coloneqq\phi_{i}(x_{i})$ and $h_{n}\coloneqq\phi_{n}(x_{n})$ evenly, splitting them into two parts each. For instance, if the original dimension of $h_{i}$ is 512, each resulting vector will have a dimension of 256. We then perform similarity and dissimilarity learning on these vectors. Specifically, we define the two parts of the split vector as the similar vectors, denoted as $h_{i}^{sim}$ and $h_{n}^{sim}$, and the dissimilar vectors, denoted as $h_{i}^{dis}$ and $h_{n}^{dis}$, obtaining $h_{i}=h_{i}^{sim}\oplus h_{i}^{dis}$ and $h_{n}=h_{n}^{sim}\oplus h_{n}^{dis}$. We enforce the dissimilar vectors to be distant from one another, while ensuring the similar vectors remain close together. This is expressed as $\mathcal{L_{S}}=\frac{1}{N}\sum_{k=1}^{N}||h_{i}^{sim}-h_{n}^{sim}||_{2}^{2},~{}\mathcal{L_{D}}=\frac{1}{N}\sum_{k=1}^{N}\left(\max\left(0,5-||h_{i}^{dis}-h_{n}^{dis}||_{2}^{2}\right)\right)_{k},$ where $\mathcal{L_{S}}$ and $\mathcal{L_{D}}$ denote similarity loss and dissimilarity loss, respectively. The total loss is $\mathcal{L_{R}}+\mathcal{L_{D}}+\mathcal{L_{S}}$.
@@
After training the encoder-decoder model, we lock the parameters of both encoders $\phi_{i}$ and $\phi_{n}$. Then, we apply $\phi_{i}$ to $\mathcal{S}_{i}$, $\phi_{n}$ to $\mathcal{S}_{n}$, and obtain an embedding of interpolatable features $\mathcal{H}_{i}:=\phi_{i}(\mathcal{S}_{i})$ and an embedding of non-interpolatable features $\mathcal{H}_{n}:=\phi_{n}(\mathcal{S}_{n})$. Note that $\mathcal{H}_{i}=concat(\mathcal{H}_{i}^{sim},\mathcal{H}_{i}^{dis})$, where $\mathcal{H}_{i}^{sim}$ is the invariant embedding of interpolatable features containing all $h_{i}^{sim}$, and $\mathcal{H}_{i}^{dis}$ is the dissimilar embedding of interpolatable features containing all $h_{i}^{dis}$. $concat$ is the pairwise feature concatenation of $(h_{i}^{sim},h_{i}^{dis})$ such that $h_{i}^{sim},h_{i}^{dis}$ belong to the same $h_{i}$. Similarly, $\mathcal{H}_{n}=concat(\mathcal{H}_{n}^{sim},\mathcal{H}_{n}^{dis})$.
@@
As shown in Figure 4, we refer to $\mathcal{H}_{n}$ as the N-retrieval embedding, and $\mathcal{H}_{i}^{sim}\cup\mathcal{H}_{n}^{sim}$, which is the invariant embedding of both interpolatable and non-interpolatable features, as I-alignment embedding. Note that a one-to-one correspondence exists between $\mathcal{S}$ (the original malware feature representation set), the N-retrieval embedding (representing non-interpolatable features), and the I-alignment embedding (representing common aspects between interpolatable and non-interpolatable features). For every original malware sample in $\mathcal{S}$, there is a projection of its non-interpolatable features in the N-retrieval embedding. Similarly, the shared aspects between each sample’s interpolatable and non-interpolatable features are mapped onto the I-alignment embedding.
@@
Once the N-retrieval and I-alignment databases are obtained, we proceed to steps (iii) and (iv), which mutate a malware instance by mixing it with a similar instance using these databases. The motivations for establishing these databases, as well as using them in steps below, are presented in our insights in Section 1 and Section 3.3. Steps (iii) and (iv) are illustrated by Figure 5.
@@
Figure 6: Illustration of similarities shared among families in BODMAS. The left subfigure displays such similarities in the original static feature space $\mathcal{S}$. The middle subfigure depicts shared similarities in the N-Retrieval embedding $\mathcal{H}_{n}$. The right subfigure represents similarities between interpolatable and non-interpolatable features of similar malware when projected to the I-Alignment invariance embedding $\mathcal{H}_{i}^{sim}\cup\mathcal{H}_{n}^{sim}$.
@@
For task (a), we identify the top-k similar malware samples in $\mathcal{S}$ compared to $s$ through constructing a k-nearest-neighbor graph. We select L2 distance computed on initial features as the similarity metric for the $k$-NN graph. From the top-k similar malware samples, we randomly select one neighbor, denoted $s^{\prime}$.
@@
where $\alpha$ is the same $\alpha$ used above. We consequently select the top-k nearest neighbors of $\tilde{h}_{n}$ in $\mathcal{H}_{n}$ using L2 distance as similarity metric. The selected neighbors, denoted $\{\hat{h}_{n_{1}},\cdots,\hat{h}_{n_{k}}\}$, are regarded as candidates in the N-Retrieval embedding space representing a mixture of non-interpolatable features of $s$ and $s^{\prime}$. A more detailed motivation for performing task (c) and the following step (iv) in the embedding spaces, rather than of the original space, will be provided in Section 3.3.
@@
In doing so, we choose one set of non-interpolatable features that aligns most closely with the mixed interpolatable features in the I-alignment embedding. Recall that $\hat{h}_{n_{j}}$ has a one-on-one correspondence with some $\tilde{s}_{n}\in\mathcal{S}_{n}$ in the original feature space. As a result, we use $\tilde{s}_{n}$ as the mixture of $s$ and $s^{\prime}$ by their non-interpolatable features. We define our data augmentation method $\mathcal{A}$ on $s\in\mathcal{S}$ such that $\mathcal{A}(s)=\tilde{s}_{i}\oplus\tilde{s}_{n}$.
@@
In essence, $\mathcal{A}(s)$ is a concatenation of a direct mutation on the interpolatable features of $s$ and a real set of non-interpolatable features retrieved from another training sample. Steps (iii) and (iv) address the problem that arises when mixing non-interpolatable features from two samples, which typically results in a semantically nonsensical mixture. This issue is mitigated by selecting non-interpolatable features from real samples that are akin to the mixed instance (step (iii)) and compatible with the mutated interpolatable features (step (iv)).
@@
During the data augmentation process, establishing a one-on-one between space $\mathcal{S}$ and the hidden spaces $\mathcal{H}_{i},\mathcal{H}_{n}$ allows the model to optimize across both the original interpolatable feature space and the non-interpolatable hidden space. This results in a more realistic synthetic sample. We provide a detailed justification for our design choices in the next subsection and a intuitive walkthrough of our proposed data augmentation pipeline in Section 4. The data augmentation method is formulated in Algorithm 1 in the Appendix.
@@
Figure 7: Case study workflow. In the left subfigure, we identify several non-interpolatable features representing a blend of the two chosen samples in the embedding space. In the middle subfigure, we select the non-interpolatable features that best align with the combined interpolatable features in the embedding. In the right subfigure, we map the synthetic malware, consisting of the chosen mixed features, back to the original feature space.
3.3 Assumptions and Insights
We make three key assumptions that underpin our data augmentation approach based on the insights described in Section 1. In this subsection, we describe and validate these assumptions.
(i) We assume that same-family malware are similar in the original static feature space ($\mathcal{S}$). Consequently, we mutate each malware sample $s$ by mixing it with one of its top-k neighbors, presuming they are instances of the same family and thus semantically similar (see Section 3.2 part (iii) task (a)). We hypothesize that mixing two such samples likely yields a new sample with more realistic features than existing data augmentation techniques.
(ii) We assume that non-interpolatable features from similar malware are projected closer together within an embedding space. Thus, we expect that mixing these features’ projections in N-Retrieval embedding $\mathcal{H}_{n}$ via linear combination will neighbor other non-interpolatable features corresponding to similar functionality (see Section 3.2 part (iii) tasks (b) & (c)).
(iii) We assume that using alignment techniques, interpolatable and non-interpolatable features from the same malware are projected onto similar subspaces within an embedding space. Thus, similarity within the I-Alignment invariance embedding $\mathcal{H}_{i}^{sim}\cup\mathcal{H}_{n}^{sim}$ indicates compatibility between a pair of interpolatable and non-interpolatable features (i.e., some pairs of features may be interdependent). Accordingly, we use similarity search between interpolatable and non-interpolatable features in this embedding space to align them and maximize feature consistency within samples we synthesize (see Section 3.2 part (iv)).
Similarity analysis Figure 6 visualizes these insights using T-SNE [45] over samples in the BODMAS dataset. We measure similarity using Euclidean distance [45]. Assumption (i) is supported by the organized clusters of malware families in Figure 6-left. Numerically, we also verified that, within BODMAS-20, a subset of BODMAS detailed in Section 5.2, each malware sample shares the same family as $95.86\%$ of its top-5 nearest neighbors. For assumption (ii), Figure 6-middle shows non-interpolatable features form family-based clusters in the N-Retrieval embedding, $\mathcal{H}_{n}$. In contrast, most non-interpolatable features do not form clusters in the original feature space $\mathcal{S}_{n}$. Thus, we mix non-interpolatable features in the embedding space $\mathcal{H}_{n}$ (Equation (3)). Finally, assumption (iii) is supported by the overlapping subspaces of same-family features in the invariance embedding $\mathcal{H}_{i}^{sim}\cup\mathcal{H}_{n}^{sim}$, as illustrated in Figure 6-right. Consequently, similarity between projected interpolatable and non-interpolatable features indicates higher compatibility.
Note that our approach operates primarily in the projected embedding space. Our assumptions and similarity analysis demonstrate our manipulations yield meaningful relationships in this space, but they do not fully translate to an explanation of specific malware features, traits, or functionalities. In this regard, this Section as well as Section 4 below, provide preliminary evidence that our approach captures relevant semantic patterns, rather than as an attempt to achieve full explainability of the underlying problem domain.
4 Motivating Example